# Microphone Array Geometry Independent Multi-Talker Distant ASR: NTT System for the DASR Task of the CHiME-8 Challenge

### v1

@@ 7.4 ASR and diarization error distribution
Figure 11: Distribution of ASR and diarization errors for the different datasets of the evaluation set of our best system (DIA3 SE ASR5). The ASR errors are measured in terms of insertions (INS), deletions (DEL), and substitutions (SUB). The diarization errors are measured in terms of missed speech (MI), false alarm (FA), and confusion (CF).
Figure 11 shows the distribution of ASR and diarization errors for the different datasets obtained with our top-performing system (system 5 of Table 2). We observe that deletion errors are the main cause of errors for tcpWER and false alarms for diarization. The number of confusion errors is also much higher for CHiME-6 than for the other datasets.
False alarms dominate diarization errors, but insertions do not dominate recognition errors, suggesting a complex relationship between diarization and recognition errors. In fact, as seen in A.1.4, short-pause filling, which increases false alarm, reduces tcpWER as the ASR backends can benefit from longer segments and ignore short pauses in-between.
In addition, we should not forget that speaker confusion errors of diarization translate into deletion and insertion errors in terms of tcpWER because the speech can be recognized but is allocated to the wrong speaker. This may explain a large number of deletions for the CHiME-6 data and indicate that reducing confusion errors on this data would be important to further improve performance by, e.g., using more discriminative speaker embeddings for EEND-VC.
@@ 7.5 Comparison of NOTSOFAR-1 results with CHiME-8 task 2 systems
For the NOTSOFAR-1 data, we can compare the results of our microphone array geometry independent systems with those submitted to the CHiME-8 task 2, which allowed developing systems dedicated only to the NOTSOFAR-1 data. In addition to the fact that the domain is limited to office meetings, the task 2 participants could take advantage of the fact that all NOTSOFAR-1 recordings were performed with a microphone array with the same known array geometry.
@@
Table 3: Comparison of task 2 (geometry dependent) and task 1 (geometry independent) systems of the CHiME-8 challenge on the NOTSOFAR-1 data in terms of session-wise macro tcpWER on the evaluation set. The best performance among all systems is indicated in bold fonts. The best performance among the array geometry independent systems is underlined.
@@ 7.6 System complexity
Table 4: Comparison of RTF and tcpWER for the three proposed systems on the development set of NOTSOFAR-1.
@@
The three systems we proposed have very different computational complexity. To provide an idea of the complexity of the different systems, we calculated the real-time factor (RTF). Table 4 reports the RTF and the tcpWER of the three proposed systems, computed on one A6000 GPU on the NOTSOFAR-1 dataset.
@@
Note that the post-challenge system DIA3 has the same complexity as DIA2 since it uses the same EEND-VC module, and the model configurations of the TS-VAD models are the same; only the training data differ.
@@ 8 Conclusion
We have introduced a microphone array geometry independent multi-talker DASR system we submitted to the DASR task of CHiME-8. We have provided a detailed description of all parts of our system and performed a detailed analysis and ablation of the effect of each component.
Our best post-challenge system could greatly improve macro tcpWER performance over the baseline by up to 63 % and outperforms the top challenge system in terms of DER. Besides, on the NOTSOFAR-1 data, we achieved new state-of-the-art results for the array-geometry agnostic system, narrowing the gap to about two points with the challenge of the best array-geometry dependent system.
Our analysis has also identified important future research directions. First, we should improve speaker counting further, as small errors can lead to large performance degradation. Moreover, we should tackle the high confusion error rate on the CHiME-6 data by developing more robust speaker embedding extractor and clustering approaches. Finally, high performance comes at the price of high system and computational complexity. Simplifying our system and reducing its computational complexity while maintaining performance will also be part of our future research directions.
@@ Appendix A Ablation analysis
We perform ablation studies on the diarization, SE, and ASR modules to explain our implementation choices.
@@ A.1 Ablation experiments on the diarization module
In this section, we report DER values optimized for diarization and not for ASR. Therefore, there is a slight discrepancy between the results of Tables 1 and those in this section.
@@ A.1.1 Diairization model configuration
Table 5: DER [%] (($\downarrow$( ↓) on the development set for different configurations of the diarization system. The DER was computed with md-eval with a collar of 0.25 sec.
@@
Table 5 shows the DER for different configurations of the EEND-VC pre-diarization and the impact of pre-diarization on TS-VAD performance as discussed in Section 3. First, comparing rows 1 and 2, we see that using SE before computing speaker embeddings as explained in Section 3.1 has a great impact on the performance as it reduces the macro DER by close to six points. This indicates that by using SE, we can obtain more discriminative speaker embeddings, improving speaker embedding clustering accuracy. Using SE also impacts speaker counting performance, as shown in A.1.2. The effect of SE is consistent on all datasets.
Next, we confirm the importance of a good initialization for TS-VAD. For this, we compare results with TS-VAD using our proposed EEND-VC (row 3) and NeMo baseline (row 4) as pre-diarization. These results show that TS-VAD is much more effective and achieves greatly lower DER when using EEND-VC.
@@ A.1.2 Speaker Counting
Table 6 shows the speaker counting accuracy and error on the development set for different configurations explained in Section 3.2. System 0 presents the results from the NeMo-based baseline system, which failed to correctly estimate the number of speakers in most scenarios except for Mixer 6. It tended to overestimate the number of speakers for CHiME-6 and DiPCo, probably because it uses short segments of up to three seconds. This caused significant variations within a speaker’s voice in highly overlapped chunks, leading to an overestimation of the number of speakers. Conversely, on NOTSOFAR-1, the baseline underestimated the number of speakers because the sessions were shorter. This resulted in less reliable clustering because of the limited number of speaker embeddings available.
@@
From Table 6, we can see that our baseline (system 1) outperforms the NeMo-based baseline in terms of the macro average computed across all scenarios. The improved performance can be attributed to using longer subchunks, i.e., 15 vs. three seconds for the NeMo baseline. However, system 1 tends to underestimate the number of speakers, except for Mixer 6, probably because, in longer subchunks, errors in estimating speaker activation often result in extracting speaker embeddings that include speech from other speakers or noise segments. As a result, the variance of the speaker embeddings increases, causing different clusters to be counted as a single cluster. On the other hand, System 2 significantly improves performance, as the interfering speakers’ speech and noise are suppressed, resulting in less variation in the embeddings.
Using microphone group-wise counting (system 3) and enhanced signals (system 4) improved overall counting performance. Notably, the performance improved for the shorter sessions of NOTSOFAR-1, likely because group-wise counting helped increase the number of embeddings, thereby enhancing performance. Moreover, applying group averaging (system 5) improved further counting accuracy, achieving the best macro-average performance across all scenarios (89.6 in SCA of Macro Avg.). These findings suggest that our proposed group-wise speaker counting on embeddings from enhanced signals, followed by aggregation through group averaging—demonstrates robust performance across various situations.
@@ A.1.3 Effect of DOVER-Lap for combining channel-wise diarization results
Figure 12: Effect of DOVER-Lap for combining the channel-wise diarization results. We report DER values for four sessions of the development set (one for each dataset). The orange and blue cross-marks show the channel-wise results using EEND-VC (DIA1) and TS-VAD (DIA2), respectively. The grey and black bars show the results obtained after using DOVER-Lap.
We perform diarization for each microphone and combine the results with DOVER-Lap as explained in Section 3.3. Figure 12 shows the DER on one session of the development set for each dataset with and without TS-VAD refinement and before and after using DOVER-Lap. We observe that before applying DOVER-LAP, the DER depends greatly on the microphone used. This tendency is especially visible with the DiPCo data, where the microphones are more widely distributed compared to those in other datasets. DOVER-Lap effectively combines the results obtained with each microphone and achieves the lowest DER, which is usually lower than the best microphone output. This experiment confirms the effectiveness of using DOVER-Lap to combine the diarization output of all the microphones in a distributed microphone array setting.
@@ A.1.4 Tuning Diarization hyperparameters for ASR
The diarization post-processing is used to convert the speaker activity posteriors into utterances boundaries, as explained in Sections 3.1.4 and 3.4. In this section, we investigate the impact of the post-processing hyperparameters on DER and tcpWER, including the boundary offset duration $\theta_{\textrm{offset}}$, the short-pause filling duration $\theta_{\textrm{merge}}$, and the activity threshold $\theta_{\textrm{ts-vad}}$.
@@
Figure 13 shows the relationship between tcpWER and DER for the different tuning parameters when using DIA2 and ASR1. The three figures are identical and differ only by their labels, indicating different values of the hyper-parameters. The optimal hyperparameters are different for diarization and ASR. We observe that the configuration achieving the best tcpWER (about 23.5 %) is close to 3 points behind the configuration achieving the best DER (about 15.2 %).
@@
Looking into each parameter in more detail, we can see from Fig. 13(a) that the offset does not have a major impact, and setting it to a small value ($\theta_{\textrm{offset}}=0.0$ or $0.05$) achieves the best tcpWER.
@@
Looking at Fig. 13(b), we see that merging segments together by filling short pauses has a larger impact on tcpWER, and the best performance is achieved with a filling duration of $\theta_{\textrm{merge}}=1.5$ sec. These results suggest that the diarization system may tend to output short segments by cutting speech at every short pause. However, for ASR, short pauses may not matter much; instead, it may be beneficial to process longer segments to exploit speech context.
@@
Finally, from Fig. 13(c), we also observe that the activity threshold also plays an important role. The optimal threshold for tcpWER ($\theta_{\textrm{ts-vad}}=0.30$) tends to be lower than that for DER ($\theta_{\textrm{ts-vad}}=0.50$), which indicates that the ASR system is less affected by false alarms than missed speech.
A.2 Ablation experiments on the SE module
A.2.1 MIMO beamforming, reference microphone selection, and postfiltering
@@
We tested whether SP-MWF was more effective than R1-MWF used by the baseline. We also investigated whether the BAN postfilter was effective in conjunction with SP-MWF. The hyperparameter $\gamma$ was set to $\gamma=0$ for both SP-MWF and R1-MWF as in the baseline.
@@
Table 7 shows the tcpWER on the development set when the beamforming and postfilters are varied in the official baseline SE system with oracle diarization. To evaluate the tcpWER, we used the baseline NeMo ASR system. The number of microphones selected in the microphone subset selection module was tested for $K=M$ and $K=\lceil 0.8\times M\rceil$. As discussed in Section 4.3.6, when the BAN postfilter is applied, the performance difference between R1-MWF and SP-MWF is attributed solely to their effectiveness of reference microphone selection. Thus, by comparing lines (1) with (3) and (5) with (7) in Table 7, we observed that SP-MWF selected a more effective reference microphone than R1-MWF. This trend was evident in the CHiME-6 and DiPCo datasets. This is likely because in these datasets the microphone arrays are more widely distributed, making the selection of an effective reference microphone more important than in the other two datasets.
@@
Next, by comparing lines (3) with (4) and (7) with (8), we observed that the BAN postfilter was not effective with SP-MWF. Applying the BAN postfilter is independent of the choice of the reference microphone and only alters the frequency-wise amplitude of the separated signal. Therefore, it turned out that the scale of SP-MWF was more appropriate than the scale defined by the BAN postfilter, although the performance improvement was small.
@@
Table 7: tcpWER [%] ($\downarrow$) on the development set when varying the beamformers and postfilters in the SE system. The results were obtained with oracle diarization and the NeMo ASR system.
@@
ID #mics ($K$) Beamformer Postfilter CHiME-6 DiPCo Mixer 6 NOTSOFAR-1 Macro
@@
(5) $\lceil 0.8\times M\rceil$ R1-MWF BAN 20.01 31.52 10.87 13.64 19.01
@@ A.2.2 Microphone subset selection
We evaluate the effect of introducing $K_{\min}$, the minimum number of microphones selected in microphone subset selection (see Section 4.2.2). We also demonstrate the effectiveness of exploiting two acoustic features, EV and $C_{50}$.
Table 8 shows the tcpWER on the development set as we vary the value of $K_{\min}$ among $\{0,5,10,15,20\}$. To evaluate the tcpWER, we used the ASR system, WavLM Transducer (ASR4 in Table 10). Except for $K_{\min}$, we used our proposed full SE system with oracle diarization. Since there are only $M=10$ microphones in Mixer 6 and $M=7$ in NOTSOFAR-1, all the microphones are selected for these datasets when $K_{\min}\geq 10$, resulting in the same tcpWER (note that the tiny difference of 0.01 % in tcpWER may be caused by numerical errors). For all datasets, the result with $K_{\min}\leq 5$ is worse than that with $K_{\min}\geq 10$, particularly for Mixer 6 and NOTSOFAR-1, showing the effectiveness of the mechanism that selects at least $K_{\min}$ microphones. For CHiME-6, the result with $K_{\min}=20$ is significantly worse than that with $K_{\min}=15$. Since CHiME-6 has only 20 or 24 microphones, when $K_{\min}=20$, all or almost all of the microphones are selected, including those that may be problematic for the SE system. By setting $K_{\min}$ to 15, we can discard these problematic microphones while still selecting a sufficient number of microphones to improve the accuracy of multichannel SE methods, such as beamforming and GSS. Based on these observations, we conclude that the $K_{\min}$ mechanism is essential for CHiME-8.
Next, Table 9 shows the tcpWER on the development set when using either or both of the EV and $C_{50}$ in the proposed SE system with oracle diarization. To evaluate the tcpWER, we used the baseline NeMo ASR system. Note that the results for Mixer 6 and NOTSOFAR-1 are the same across all methods since we select all $M$ microphones due to $K_{\min}=15$. The results for CHiME-6 and DiPCo show that relying on two features is more advantageous than relying on just a single feature, clearly demonstrating the effectiveness of the proposed approach. In future works, we plan to develop a method that can handle more than two audio features and explore more sophisticated yet simple combinations of these features.

### v2

@@ 7.4 ASR and diarization error distribution
Figure 11: Distribution of ASR and diarization errors for different datasets of the evaluation set of our best system (DIA3 SE ASR5). ASR errors are measured in terms of insertions (INS), deletions (DEL), and substitutions (SUB). Diarization errors are measured in terms of missed speech (MI), false alarm (FA), and confusion (CF).
Figure 11 shows the distribution of ASR and diarization errors for the different datasets obtained with our top-performing system (system 5 in Table 2). We observe that deletion errors are the main cause of errors for tcpWER and false alarms for diarization. The number of confusion errors is also much higher for CHiME-6 than for the other datasets.
False alarms dominate diarization errors, but insertions do not dominate recognition errors, suggesting a complex relationship between diarization and recognition errors. In fact, as discussed in A.1.4, short-pause filling, which increases false alarm, reduces tcpWER as the ASR backends can benefit from longer segments and ignore short pauses in-between.
In addition, we should not forget that speaker confusion errors of diarization translate into deletion and insertion errors in terms of tcpWER because the speech can be recognized but is allocated to the wrong speaker. This may explain the large number of deletions for the CHiME-6 data and suggests that reducing confusion errors on this data would be important to further improve performance by, e.g., using more discriminative speaker embeddings for EEND-VC.
@@ 7.5 Comparison of NOTSOFAR-1 results with CHiME-8 task 2 systems
For the NOTSOFAR-1 data, we can compare the results of our microphone array geometry-independent systems with those submitted to the CHiME-8 task 2, which allowed developing systems dedicated only to the NOTSOFAR-1 data. In addition to the fact that the domain is limited to office meetings, the task 2 participants could take advantage of the fact that all NOTSOFAR-1 recordings were performed with a microphone array having the same known array geometry.
@@
Table 3: Comparison of task 2 (geometry-dependent) and task 1 (geometry-independent) systems of the CHiME-8 challenge on the NOTSOFAR-1 data in terms of session-wise macro tcpWER on the evaluation set. The best performance among all systems is indicated in bold fonts. The best performance among the array geometry-independent systems is underlined.
@@ 7.6 System complexity
Table 4: Comparison of RTF and tcpWER for three proposed systems on development set of NOTSOFAR-1.
@@
The three systems we proposed have very different computational complexities. To provide an idea of the complexities, we calculated the real-time factor (RTF). Table 4 reports the RTF and the tcpWER of the three proposed systems, computed on one A6000 GPU on the NOTSOFAR-1 dataset.
@@
Note also that the post-challenge system DIA3 has the same complexity as DIA2 since it uses the same EEND-VC module, and the model configurations of the TS-VAD models are the same; only the training data differ.
@@ 8 Conclusion
We have introduced the microphone array geometry-independent multi-talker DASR system we submitted to the DASR task of CHiME-8. We have provided a detailed description of all parts of our system and performed a detailed analysis and ablation study of the effect of each component.
Our best post-challenge system significantly improved the macro tcpWER performance over the baseline by up to 63 % and outperformed the top challenge system in terms of DER. Moreover, on the NOTSOFAR-1 data, we achieved new state-of-the-art results for the array-geometry agnostic system, narrowing the gap to about two points behind the best array-geometry-dependent system in the challenge.
Our analysis has also identified several important future research directions. First, we should improve speaker counting further, as small errors can lead to large performance degradation. Moreover, we should tackle the high confusion error rate on the CHiME-6 data by developing more robust speaker embedding extractor and clustering approaches. Finally, high performance comes at the price of high system and computational complexity. Simplifying our system and reducing its computational complexity while maintaining performance will also be part of our future research directions.
@@ Appendix A Ablation analysis
We performed ablation studies on the diarization, SE, and ASR modules to clarify the effect of our implementation choices.
@@ A.1 Ablation experiments on the diarization module
In this section, we report the DER values optimized for diarization and not for ASR. Therefore, there is a slight discrepancy between the results in Tables 1 and those in this section.
@@ A.1.1 Diairization model configuration
Table 5: DER [%] (($\downarrow$( ↓) on the development set for different configurations of the diarization system. The DER was computed using md-eval with a collar of 0.25 sec.
@@
Table 5 shows the DER for different configurations of the EEND-VC pre-diarization and the impact of pre-diarization on TS-VAD performance as discussed in Section 3. First, comparing rows 1 and 2, we see that using SE before computing speaker embeddings as explained in Section 3.1 has a significant impact on the performance as it reduces the macro DER by close to six points. This indicates that by using SE, we can obtain more discriminative speaker embeddings, thereby improving the accuracy of the speaker embedding clustering. Using SE also impacts the speaker counting performance, as shown in A.1.2. The effect of SE is consistent across all datasets.
Next, we confirm the importance of a good initialization for TS-VAD by comparing the results with TS-VAD using our proposed EEND-VC (row 3) and NeMo baseline (row 4) as pre-diarization. These results demonstrate that TS-VAD is much more effective and achieves a significantly lower DER when using EEND-VC.
@@ A.1.2 Speaker Counting
Table 6 shows the speaker counting accuracy and error on the development set for the different configurations explained in Section 3.2. System 0 presents the results from the NeMo-based baseline system, which failed to correctly estimate the number of speakers in most scenarios except for Mixer 6. It tended to overestimate the number of speakers for CHiME-6 and DiPCo, probably because it uses short segments of up to three seconds. This caused significant variations within a speaker’s voice in highly overlapped chunks, leading to an overestimation of the number of speakers. Conversely, on NOTSOFAR-1, the baseline underestimated the number of speakers because the sessions were shorter. This resulted in less reliable clustering because of the limited number of speaker embeddings available.
@@
In Table 6, we can see that our baseline (system 1) outperforms the NeMo-based baseline in terms of the macro average computed across all scenarios. The improved performance can be attributed to using longer subchunks, i.e., 15 vs. three seconds for the NeMo baseline. However, system 1 tends to underestimate the number of speakers (except for Mixer 6), probably because, in longer subchunks, errors in estimating speaker activation often result in extracting speaker embeddings that include speech from other speakers or noise segments. As a result, the variance of the speaker embeddings increases, causing different clusters to be counted as a single cluster. In contrast, system 2 significantly improves the performance, as the interfering speakers’ speech and noise are suppressed, resulting in less variation in the embeddings.
Using microphone group-wise counting (system 3) and enhanced signals (system 4) improved the overall counting performance. Notably, the performance improved for the shorter sessions of NOTSOFAR-1, likely because group-wise counting helped increase the number of embeddings, thereby enhancing performance. Moreover, applying group averaging (system 5) further improved the counting accuracy, achieving the best macro-average performance across all scenarios (89.6 in SCA of Macro Avg.). These findings suggest that our proposed group-wise speaker counting on embeddings from enhanced signals, followed by aggregation through group averaging, demonstrates a robust performance across various situations. The overall SCA in NOTSOFAR-1 was relatively lower than in the other scenarios, possibly because NOTSOFAR-1 is a unique scenario involving more speakers (up to eight) in short sessions (averaging six minutes) compared to the other scenarios [9]. This discrepancy may cause the hyper-parameters, including chunk and subchunk lengths, to overfit the other three scenarios. Note that although the SCA is low, the SCE on NOTSOFAR-1 remains reasonable (0.51), and the relative speaker error is moderate considering the higher number of speakers in the sessions of this dataset.
@@ A.1.3 Effect of DOVER-Lap for combining channel-wise diarization results
Figure 12: Effect of DOVER-Lap for combining the channel-wise diarization results. We report DER values for four sessions of the development set (one for each dataset). Orange and blue cross-marks show the channel-wise results using EEND-VC (DIA1) and TS-VAD (DIA2), respectively. Gray and black bars show the results obtained after using DOVER-Lap.
We perform diarization for each microphone and combine the results with DOVER-Lap as explained in Section 3.3. Figure 12 shows the DER on one session of the development set for each dataset with and without TS-VAD refinement and before and after using DOVER-Lap. We observe that before applying DOVER-LAP, the DER depends greatly on the microphone used. This tendency is especially visible with the DiPCo data, where the microphones are more widely distributed compared to those in other datasets. DOVER-Lap effectively combines the results obtained with each microphone and achieves the lowest DER, which is usually lower than the best microphone output. These findings confirm the effectiveness of using DOVER-Lap to combine the diarization output of all the microphones in a distributed microphone array setting.
@@ A.1.4 Tuning diarization hyperparameters for ASR
Diarization post-processing is utilized to convert the speaker activity posteriors into utterances boundaries, as explained in Sections 3.1.4 and 3.4. In this section, we investigate the impact of the post-processing hyperparameters on DER and tcpWER, including the boundary offset duration $\theta_{\textrm{offset}}$, the short-pause filling duration $\theta_{\textrm{merge}}$, and the activity threshold $\theta_{\textrm{ts-vad}}$.
@@
Figure 13 shows the relationship between tcpWER and DER for the different tuning parameters when using DIA2 and ASR1. The three graphs are identical and differ only by their labels, indicating different values of the hyper-parameters. The optimal hyperparameters are different for diarization and ASR. We observe that the configuration achieving the best tcpWER (about 23.5 %) is close to 3 points behind the configuration achieving the best DER (about 15.2 %).
@@
Looking into each parameter in more detail, we can see from Figure 13(a) that the offset does not have a major impact, and setting it to a small value ($\theta_{\textrm{offset}}=0.0$ or $0.05$) achieves the best tcpWER.
@@
Looking at Figure 13(b), we see that merging segments together by filling short pauses has a larger impact on tcpWER, and the best performance is achieved with a filling duration of $\theta_{\textrm{merge}}=1.5$ sec. These results suggest that the diarization system may tend to output short segments by cutting speech at every short pause. However, for ASR, short pauses may not matter much; instead, it may be beneficial to process longer segments to exploit speech context.
@@
Finally, from Figure 13(c), we also observe that the activity threshold plays an important role. The optimal threshold for tcpWER ($\theta_{\textrm{ts-vad}}=0.30$) tends to be lower than that for DER ($\theta_{\textrm{ts-vad}}=0.50$), which indicates that the ASR system is less affected by false alarms than missed speech.
A.2 Ablation experiments on SE module
A.2.1 MIMO beamforming, reference microphone selection, and post-filtering
@@
We tested whether SP-MWF was more effective than the R1-MWF utilized by the baseline. We also investigated whether the BAN post-filter was effective in conjunction with SP-MWF. The hyperparameter $\gamma$ was set to $\gamma=0$ for both SP-MWF and R1-MWF as in the baseline.
@@
Table 7 shows the tcpWER on the development set when the beamforming and post-filters are varied in the official baseline SE system with Oracle diarization. To evaluate the tcpWER, we used the baseline NeMo ASR system. The number of microphones selected in the microphone subset selection module was tested for $K=M$ and $K=\lceil 0.8\cdot M\rceil$. As discussed in Section 4.3.6, when the BAN post-filter is applied, the performance difference between R1-MWF and SP-MWF is attributed solely to their effectiveness of the reference microphone selection. Thus, by comparing lines (1) with (3) and (5) with (7) in Table 7, we observed that SP-MWF selected a more effective reference microphone than R1-MWF. This trend was only evident in the CHiME-6 and DiPCo datasets, likely because in these datasets the microphone arrays in these datasets are more widely distributed, making the selection of an effective reference microphone more important than in the other two datasets.
@@
Next, by comparing lines (3) with (4) and (7) with (8), we observed that the BAN post-filter was not effective with SP-MWF. Applying the BAN post-filter is independent of the choice of the reference microphone and only alters the frequency-wise amplitude of the separated signal. Therefore, it turned out that the scale of SP-MWF was more appropriate than the scale defined by the BAN post-filter, although the performance improvement was small.
@@
Table 7: tcpWER [%] ($\downarrow$) on the development set when varying the beamformers and post-filters in the SE system. The results were obtained with Oracle diarization and the NeMo ASR system.
@@
ID #mics ($K$) Beamformer post-filter CHiME-6 DiPCo Mixer 6 NOTSOFAR-1 Macro
@@
(5) $\lceil 0.8\cdot M\rceil$ R1-MWF BAN 20.01 31.52 10.87 13.64 19.01
@@ A.2.2 Microphone subset selection
We evaluate the effect of introducing $K_{\min}$, the minimum number of microphones selected in microphone subset selection (see Section 4.2.2). We also investigate the effectiveness of exploiting two acoustic features, EV and $C_{50}$.
Table 8 shows the tcpWER on the development set as we vary the value of $K_{\min}$ among $\{0,5,10,15,20\}$. To evaluate the tcpWER, we used the ASR system, WavLM Transducer (ASR4 in Table 10). Except for $K_{\min}$, we used our proposed full SE system with Oracle diarization. Since there are only $M=10$ microphones in Mixer 6 and $M=7$ in NOTSOFAR-1, all the microphones are selected for these datasets when $K_{\min}\geq 10$, resulting in the same tcpWER (note that the tiny difference of 0.01 % in tcpWER may be caused by numerical errors). For all datasets, the result with $K_{\min}\leq 5$ is worse than that with $K_{\min}\geq 10$, particularly for Mixer 6 and NOTSOFAR-1, demonstrating the effectiveness of the mechanism that selects at least $K_{\min}$ microphones. For CHiME-6, the result with $K_{\min}=20$ is significantly worse than that with $K_{\min}=15$. Since CHiME-6 has only 20 or 24 microphones, when $K_{\min}=20$, all or almost all of the microphones are selected, including those that may be problematic for the SE system. By setting $K_{\min}$ to 15, we can discard these problematic microphones while still selecting a sufficient number of microphones to improve the accuracy of multichannel SE methods, such as beamforming and GSS. On the basis of these observations, we conclude that the $K_{\min}$ mechanism is essential for CHiME-8.
Next, Table 9 shows the tcpWER on the development set when using either or both of the EV and $C_{50}$ in the proposed SE system with Oracle diarization. To evaluate the tcpWER, we used the baseline NeMo ASR system. Note that the results for Mixer 6 and NOTSOFAR-1 are the same across all methods since we select all $M$ microphones due to $K_{\min}=15$. The results for CHiME-6 and DiPCo show that relying on two features is more advantageous than relying on just a single feature, clearly demonstrating the effectiveness of the proposed approach. In future work, we plan to develop a method that can handle more than two audio features and explore more sophisticated yet simple combinations of these features.